OpenStudy (anonymous):

find the lenght of c if angle A=62 angle B=39 and side b =15

OpenStudy (anonymous):

triangle? use law of sines (sinB)/b=(sinC)/c C=180-A-B

OpenStudy (anonymous):

\[c=\frac{b \sin C}{\sin B}\]

OpenStudy (anonymous):

the choices are A 10.7 B 23.4 C 27.4 D 13.3

OpenStudy (anonymous):

if u use the formula that dockworker gave then u should have 10.7,

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

its not 10.7, remember to use degree mode in your calculator